OpenStudy (anonymous):

Is there a combination of 15 cent and 33 cent stamps that can make $1.77? If so, find all possible ways of doing this. If not, explain clearly why not.

ganeshie8 (ganeshie8):

you want to solve \[15a + 33b = 177\] in integers

OpenStudy (anonymous):

would i have to cancel one out by doing 15a(-15a)?

OpenStudy (anonymous):

oh! would i have to do them from 1 all the way up until i get to 177 and stop when i get there? like 15(1)+33(1)=177?

ganeshie8 (ganeshie8):

That works! but there are more efficinet ways... familiar with finding gcd using euclid algorithm ?

OpenStudy (anonymous):

no, i haven't learned the euclid algorithm :o

ganeshie8 (ganeshie8):

Okay, lets not use it then. First of all notice that this equation has solutions in integers because both sides are multiples of \(3\). May be lets divide that

ganeshie8 (ganeshie8):

\[15a + 33b = 177\] divide 3 through out \[5a + 11b = 59\]

ganeshie8 (ganeshie8):

Clearly \(b\) cannot be more than \(5\) (why ?) consider five cases \(b = 1,2,3,4,5\) and see which value(s) of \(b\) give you integer values for \(a\)

OpenStudy (anonymous):

why can't b be more than 5?

OpenStudy (anonymous):

isn't 11 already greater than 5?

ganeshie8 (ganeshie8):

you want 5a+11b to equal 59 when b is greater than 5, the left hand side will be greater than 59. so NO solutions.

OpenStudy (anonymous):

oh okay! so then would my equation be: 5(1)+11(1)=59, then all the way up to 5?

ganeshie8 (ganeshie8):

consider 5 cases

ganeshie8 (ganeshie8):

case 1 : b=1 \(5a + 11(1) = 59\) solve \(a\)

OpenStudy (anonymous):

i got 9.6...

ganeshie8 (ganeshie8):

which is not an integer, so case1 wont work

ganeshie8 (ganeshie8):

work out case2

ganeshie8 (ganeshie8):

case 2 : b=2 \(5a + 11(2) = 59\) solve \(a\)

OpenStudy (anonymous):

7.4, so that doesnt work..

ganeshie8 (ganeshie8):

Right, work remaining 3 cases also

OpenStudy (anonymous):

only case 4 works, a=3

ganeshie8 (ganeshie8):

does that mean the only solution is a=3, b=4 ?

OpenStudy (anonymous):

yes, because they are equal right?

OpenStudy (anonymous):

or do i need to create another equation?

ganeshie8 (ganeshie8):

Yep!

ganeshie8 (ganeshie8):

No, we're done

OpenStudy (anonymous):

awesome, thank you so much!!

OpenStudy (anonymous):

so now if i wanted to make $1.77 using only 15 and 35 cent stamps how could i find all the possible ways for that?

OpenStudy (anonymous):

could i still use a = 3 and b = 4?

ganeshie8 (ganeshie8):

yes we figured out that the only possible combination is by taking three 15 cent stamps and four 35 cent stamps

ganeshie8 (ganeshie8):

lets double check maybe..

ganeshie8 (ganeshie8):

3(0.15) + 4(0.35) = ?

OpenStudy (anonymous):

wait i got 1.85!

ganeshie8 (ganeshie8):

3(0.15) + 4(0.33) = ?

OpenStudy (anonymous):

let me check my math again!

ganeshie8 (ganeshie8):

there was a typo, its 33 cent stamps right ?

OpenStudy (anonymous):

yes!

ganeshie8 (ganeshie8):

3(0.15) + 4(0.33) = ?

OpenStudy (anonymous):

okay so thats 1.77! awesome, thank you so much!! you definitely made it a lot easier to understand a problem that i was over thinking! how did you know to use the integers and create the equation 15a+33b=177 though?

ganeshie8 (ganeshie8):

we want to find a, b such that : 0.15a + 0.33b = 1.77

ganeshie8 (ganeshie8):

since working with integers is easy, i have simply multiplied 100 both sides :)

OpenStudy (anonymous):

oh i see! thanks!! :)

ganeshie8 (ganeshie8):

yw:)